# Probabilistic Time Integration for Semi-explicit PDAEs

### v1

@@ 5. Calibrating Forward Uncertainty Propagation
This section aims to discuss one potential approach for choosing $\sigma$ when employing randomized time integrators to solve constrained parabolic systems. Let us consider once more the constrained FitzHugh–Nagumo model from Examples 1.1 and 2.5. Through this example, we can observe that the scale parameter $\sigma$ controls the apparent uncertainty in the solver; see Figure 1.2. While the rate of contraction is governed by the underlying deterministic method, this affects the error magnitude induced by the probabilistic numerical method. To overcome this drawback, we follow [CGS+17] and calibrate $\sigma$ to match the error magnitude given by classical error indicators. Let us denote the approximation given by the probabilistic solver by $U_{\tau,\sigma}$ and the corresponding deterministic approximation by $u_{\tau,0}$. The simplest error indicator might be the difference between solutions obtained with different step sizes, i.e., $E^{n}\coloneqq u^{n}_{\tau,0}-u^{n}_{2\tau,0}$.
@@
In line with [CGS+17, Sect. 3], we study the probability distribution $\pi(\sigma)$, which achieves maximum likelihood when the desired matching occurs. This estimation of scale matching involves the comparison of a Gaussian approximation $\tilde{\mu}^{n}_{\tau,\sigma}=\mathcal{N}(\mathbb{E}(U^{n}_{\tau,\sigma}),\mathbb{V}(U^{n}_{\tau,\sigma}))$ and the natural Gaussian measure represented as $\nu^{n}_{\tau,0}=\mathcal{N}(u^{n}_{\tau,0},(E^{n})^{2})$ at each time step. Using the Bhattacharyya distance [Kai67] to measure the disparity between these two normal distributions, the probability distribution $\pi(\sigma)$ can be constructed by
$\pi(\sigma)=\prod_{n}\exp\left(-d(\tilde{\mu}^{n}_{\tau,\sigma},\nu^{n}_{\tau,0})\right).$
We use the following formulae
$d(\tilde{\mu}^{n}_{\tau,\sigma},\nu^{n}_{\tau,0})=\frac{1}{4}\frac{\left(\mathbb{E}(U^{n}_{\tau,\sigma})-u^{n}_{\tau,0}\right)^{2}}{\mathbb{V}(U^{n}_{\tau,\sigma})+(E^{n})^{2}}+\frac{1}{4}\ln\left(\frac{1}{4}\left(\frac{\mathbb{V}(U^{n}_{\tau,\sigma})}{(E^{n})^{2}}+\frac{(E^{n})^{2}}{\mathbb{V}(U^{n}_{\tau,\sigma})}+2\right)\right)$
for the Bhattacharyya distance when applied to the case of two univariate normal distributions [LB12]. This effectively assigns a penalty that decreases exponentially as the distance between the distributions increases.
@@
For the choice $\sigma^{*}=\arg\max\pi(\sigma)$, we depict the true trajectories of the $V$ species of the constrained FitzHugh–Nagumo model and $100$ realizations from the probabilistic implicit Euler method in Figure  5.1. Therein, we use various step sizes and a fixed noise scale $\sigma=0.5$ for comparison. The plots clearly show that the random realisations from the measure associated with the calibrated probabilistic implicit Euler solver contract towards the true solution faster than those obtained by probabilistic implicit Euler solver with a fixed $\sigma$. In Figure 5.2 we present a comparison of the observed variations and their mean in the calibrated probabilistic implicit Euler method and the error indicator. This comparison demonstrates a good agreement in the marginal variances.
@@ 6. Conclusion
In this paper, we have introduced and analyzed a class of probabilistic time integrators of first and second order for the numerical solution of semi-linear parabolic equations with linear constraints. Such methods incorporate random perturbations, which do not affect the constraint manifold, in order to quantify the uncertainty induced by the time integration scheme. In terms of convergence, we have studied the mean square error and showed that a balanced inclusion of perturbations does not reduce the convergence order known from the deterministic setting. More precisely, the overall order equals the minimum of the (deterministic) convergence order and the order of the introduced statistical error. All results are validated through numerical experiments.

### v3

@@ 5. Calibrating Forward Uncertainty Propagation
This section aims to discuss one potential approach for choosing $\sigma$ when employing randomized time integrators to solve constrained parabolic systems. Let us consider once more the constrained FitzHugh–Nagumo model from Examples 1.1 and 2.5. Through this example, we can observe that the scale parameter $\sigma$ controls the apparent uncertainty in the solver; see Figure 1.2. While the random draws contract towards the true solution, the scale parameter $\sigma$ affects the error magnitude introduced by the probabilistic numerical method. To overcome this drawback, we follow [CGS+17] and calibrate $\sigma$ to match the error magnitude given by classical error indicators. Let us denote the approximation given by the probabilistic solver (with step size $\tau$) by $U^{n}_{\tau,\sigma}$ and the corresponding deterministic approximation by $u^{n}_{\tau,0}$. The simplest error indicator might be the difference between solutions obtained with different step sizes, i.e., $E^{n}\coloneqq u^{n}_{\tau,0}-u^{2n}_{\tau/2,0}$.
@@
In line with [CGS+17, Sect. 3], we study a probability distribution denoted by $\pi(\sigma)$, which achieves maximum likelihood when the desired matching occurs. This estimation of scale matching involves the comparison of a Gaussian approximation $\tilde{\mu}^{n}_{\tau,\sigma}=\mathcal{N}(\mathbb{E}(U^{n}_{\tau,\sigma}),\mathbb{V}(U^{n}_{\tau,\sigma}))$ and the natural Gaussian measure represented as $\nu^{n}_{\tau,0}=\mathcal{N}(u^{n}_{\tau,0},(E^{n})^{2})$ at each time step. Here, $\mathbb{V}(U^{n}_{\tau,\sigma})$ denotes the variance of the vector-valued random variables $U^{n}_{\tau,\sigma}$. Using the Bhattacharyya distance [Kai67] to measure the disparity between these two normal distributions, the probability distribution $\pi(\sigma)$ can be constructed by
$\pi(\sigma)\propto\prod_{n}\exp\left(-d(\tilde{\mu}^{n}_{\tau,\sigma},\nu^{n}_{\tau,0})\right).$
We use the following (component-wise) formulae
(5.1) $d(\tilde{\mu}^{n}_{\tau,\sigma},\nu^{n}_{\tau,0})=\frac{1}{4}\frac{\left(\mathbb{E}(U^{n}_{\tau,\sigma})-u^{n}_{\tau,0}\right)^{2}}{\mathbb{V}(U^{n}_{\tau,\sigma})+(E^{n})^{2}}+\frac{1}{4}\ln\left(\frac{1}{4}\left(\frac{\mathbb{V}(U^{n}_{\tau,\sigma})}{(E^{n})^{2}}+\frac{(E^{n})^{2}}{\mathbb{V}(U^{n}_{\tau,\sigma})}+2\right)\right)$
for the Bhattacharyya distance when applied to the case of two univariate normal distributions [LB12]. Given a randomized solution, calculating the mean and variance is straightforward from their definitions [CB24]. Further note that equation (5.1) should be understood component-wise.
We seek the value of $\sigma$ that maximizes the function $\pi(\sigma)$. To achieve this, we employ the MATLAB optimization algorithm fminsearch. By minimizing the negative of $\pi(\sigma)$, we reformulate the problem as
$\sigma^{*}=\arg\min_{\sigma}-\pi(\sigma).$
@@
By this choice, we depict the true trajectories of the $V$ species of the constrained FitzHugh–Nagumo model and $100$ realizations from the probabilistic implicit Euler method in Figure  5.1. Therein, we use various step sizes and a fixed noise scale $\sigma=0.5$ for comparison. The plots clearly show that the random realizations from the measure associated with the calibrated probabilistic implicit Euler solver contract towards the true solution faster than those obtained by probabilistic implicit Euler solver with a fixed $\sigma$. In Figure 5.2 we present a comparison of the observed marginal variations and their mean in the calibrated probabilistic implicit Euler method and the error indicator. This comparison demonstrates a good agreement in the marginal variances. Algorithm 3 describes the process for calibrating the noise scale $\sigma$ in probabilistic numerical methods.
@@
Algorithm 3 Calibration of noise scale $\sigma$
1:Input: step size $\tau$, consistent initial data $U^{0}=u^{0}\in\mathcal{V}$, right-hand sides $f$, $g$, initial scaled noise $\sigma$, number of realizations $M$
2:for $n=0$ to $N-1$ do
3:     compute deterministic solutions $u^{n}_{\tau,0}$ and $u^{2n}_{\tau/2,0}$
4:     compute error indicator $E^{n}\coloneqq u^{n}_{\tau,0}-u^{2n}_{\tau/2,0}$
5:     compute random solution $U^{n}_{\tau,\sigma}$ using a probabilistic solver with $M$ realizations
6:     compute the mean $\mathbb{E}(U^{n}_{\tau,\sigma})$ and the variance $\mathbb{V}(U^{n}_{\tau,\sigma})$ of the random solutions
7:     compute the Bhattacharyya distance $d(\tilde{\mu}^{n}_{\tau,\sigma},\nu^{n}_{\tau,0})$ as in (5.1)
8:     set $\sigma=\arg\min_{\sigma}-\pi(\sigma)$
9:end for
@@ 6. Conclusion
In this paper, we have introduced and analyzed a class of probabilistic time integrators of first and second order for the numerical solution of semi-linear parabolic equations with linear constraints. Such methods incorporate random perturbations, which do not affect the constraint manifold, in order to quantify the uncertainty induced by the time integration scheme. In terms of convergence, we have studied the mean square error and showed that a balanced inclusion of perturbations does not reduce the convergence order known from the deterministic setting. More precisely, the overall order equals the minimum of the deterministic convergence order and the order of the introduced statistical error. All results are validated through numerical experiments.